# Simulations of Ellipsoidal Primordial Black Hole Formation

### v2

@@ 4.1 Dynamical evolution of the non-spherical gravitational collapse
We analyze two specific cases to examine the dynamics of gravitational collapse. In the first case, the fluctuation exceeds the critical threshold ($\mu_{\rm t}>\mu_{\rm c}(e,p)$), leading to black hole formation. In the second case, the fluctuation remains below the threshold ($\mu_{\rm t}<\mu_{\rm c}(e,p)$), resulting in the dispersion of the fluctuation. To study the dynamical evolution of the fluctuation, we plot the ratio of $\rho/\rho_{b}(t)$ and the 4-velocity of the fluid $U^{i}$. We focus on the case of radiation-dominated Universe $w=1/3$ and refer the reader to the appendix C for the $w=1/10$ case. We found the qualitative behaviour to be the same for both $w=1/3$ and $1/10$ except that the collapsing time is longer for $w=1/10$, which is consistent with the case of spherical simulations [29].
@@
In Figs. 4 and 5, we plot the velocity $U^{x}$ and $U^{z}$ on the $x$-$z$ plane, respectively, where we observe a highly non-spherical distribution. From the velocity plots of $U^{x}$ and $U^{z}$, we observe that the collapse initially progresses slightly faster along the $z$-axis compared to the $x$-axis, as indicated by the higher collapse velocity in $U^{z}$ (see the panels of $t/t_{H}\approx 0.04$). However, presumably because of the larger pressure gradient along the $z$-axis, the contraction along $x$-axis overtakes the contraction along $z$ axis (see the panels of $t/t_{H}\approx 40.04$ and compare the values of $U^{x}$ and $U^{z}$). Subsequently, the initial shape transitions from horizontally long to vertically long while preserving the ellipsoidal shape and reducing the size of the overdensity region (see the panels of $t/t_{H}=90.04$ and those after that in Fig. 3). At very late times, the shape becomes nearly spherical.
@@
Fig. 6 shows the the energy density, lapse function, and fluid velocity in the $x$ components on the $x$-axis. Similar behaviors are found for $y$ and $z$ axes, and we do not display them. We observe a continuous decrease in the lapse function at the centre, and we infer the formation of an apparent horizon.
@@
Finally, Fig. 11 shows the time evolution of the lapse function at the centre $\alpha(\vec{r}=0)$ for different configurations. The figure highlights the detailed bouncing behaviour of the lapse function, which indicates apparent horizon formation when $\alpha(\vec{r}=0)$ continuously decreases and approaches to zero. We can observe that the time scale of the gravitational collapse is longer for a softer equation of state (smaller $w$). The top panels display cases with the fixed amplitude $\mu_{\rm t}$. Generally, we observe that non-spherical effects ($e,p\neq 0$) tend to slow down the collapse and can significantly increase the collapse time. The bottom panels show different configurations with varying amplitudes $\mu$ while fixing $p=0$. Here, we find that as $\mu$ increases beyond the critical value for spherical collapse, the deviation from sphericity required to prevent collapse also becomes larger, showing that non-spherical effects tend to prevent black hole formation. By running multiple simulations and analyzing the bouncing behaviour of the lapse function, we identified the threshold values, which will be discussed in detail in the next section.
@@ 4.3 Critical ellipticity as a function of the initial amplitude $\mu$
Figure 13: The critical value $e_{\rm c}$ with $p=0$ is depicted as a function of $(\mu-\mu_{\rm c,sp})/\mu_{\rm c,sp}$ together with the error bars given by the intervals between the parameter values of $e$ for the corresponding two cases of the black hole formation and dissipation. The dashed magenta lines correspond to the fitting to the equation $\tilde{e}_{c}(\mu)=\mathcal{K}_{e}[(\mu-\mu_{\rm c,sp})/\mu_{\rm c,sp}]^{\gamma_{e}}$, whose values $\mathcal{K}_{e}$ and $\gamma_{e}$ are shown in the plot for the case $w=1/10$ (blue) and $w=1/3$ (red).
In Fig. 13, we present the results of our simulations. Our numerical results closely follow a power-law behavior, described by:
@@
Our numerical results in Fig. 13 demonstrate that non-spherical effects make the gravitational collapse of fluctuations harder compared to the spherical case. Moreover, due to the behaviour of Eq.(4.2), even a slight deviation from sphericity can prevent a fluctuation from collapsing into a black hole if those fluctuations have an amplitude near the threshold in the critical regime. Let us evaluate, due to the nonspherical effects, how large fraction is excluded from the total number of black holes which is supposed to be formed without taking into account the non-sphericity. The remaining fraction can be estimated by integrating the probability distribution of $e$ and $p$ in the region enclosed by $p=\pm e$ and $p=\tilde{p}_{\rm c}(e)$ with a fixed value of $\mu$. For simplicity, we approximate the line $p=\tilde{p}_{\rm c}(e)$ by $e=\tilde{e}_{\rm c}(\mu)$ of a vertical line in $e$-$p$ plane and integrate the probability distribution $\mathcal{P}_{e,p}$ over the region $\mathcal{R}$ defined by
@@
[…]
On the other hand, if we focus on the critical scaling regime, in which the amplitude is very close to $\mu_{\rm c,sp}$, the number of configurations collapsing into black holes can be significantly reduced. For instance, for $(\mu-\mu_{\rm c,sp})/\mu_{\rm c,sp}=10^{-3}$, most of the configurations are prevented from collapsing into black holes although it slightly depends on $w$ and $\nu$. Therefore the non-spherical effects can be highly significant in the critical regime, preventing a large fraction of configurations from collapsing into black holes. Nevertheless, we report in the accompanying letter [85] that, in terms of the PBH mass function, the effects of the ellipticity may not have a significant impact even on the power-law small mass tail originates from the critical behavior.
[…]
@@
The fact that we do not observe a significant impact of non-spherical effects on gravitational collapse for $w=1/10$ compared to $w=1/3$ (as shown in Fig. 13) could be attributed to the following reason. First, we should note that the collapsing dynamics of the ellipsoidal system is different from the dust case and the well-known instability shown by Lin-Mestel-Shu [86] does not apply. Here let us try to understand our results following the Jeans criterion, which states that the system is unstable against gravitational collapse if the free-fall time scale is shorter than the sound wave crossing time scale. In our setting, at least for a relatively small ellipticity, the free-fall time scale $\sim 1/\sqrt{\rho}$ would be still relevant, and it does not change much due to the small ellipticity $e$ because the product $\lambda_{1}\lambda_{2}\lambda_{3}$ is conserved at the linear order of $e$ with $p=0$. On the other hand, the sound wave crossing time scale is expected to be shorter because the sound wave along the short axis may propagate through the system in a shorter time. Therefore the gravitational collapse is expected to be harder. Although, with a given value of the initial amplitude $\mu$, this effect is expected to be more pronounced for larger $w$, the relative effect may be comparable since the threshold amplitude is smaller for a smaller value of $w$.
Once non-spherical effects become dominant impeding factors against the gravitational collapse as $w\rightarrow 0$, the Jeans criterion does not apply, and we need to consider different criteria. Non-sphericities are expected to grow during collapse [87, 86]. This could lead to complex dynamics associated with the deformation [88] and rotation of over-densities [32], and velocity dispersion potentially play a significant role [89]. Further investigation is needed to clarify this aspect.
@@ 5 Conclusions
We have also examined how the threshold for black hole formation depends on the parameters $e$ and $p$. When we fixed the amplitude as a typical value $\mu_{\rm t}$, we found that the curve $p=p_{\rm c}(e)$ which describes the boundary of the region of black hole formation on the $e$-$p$ plane fits well a superellipse curve (described by Eq.(4.1)) characterized by an exponent $n^{\pm}$, where the sign $\pm$ denotes the two branches for the region $p\gtrless 0$. It would be interesting to investigate that functional form in terms of $\mu$, and its profile dependence for developing an analytical framework similar to spherically symmetric cases [28, 29] or including non-Gaussianities [67].
@@ Appendix A Analytical formulas for the mean values of $e$ and $p$
In the large $\xi$ limit, the probability distribution Eq.(2.11) can be indeed approximated by a Gaussian and takes the simpler form:
@@ Appendix B High peaks for PBH formation assuming spherical symmetry
In this appendix, under the assumption of spherical symmetry, we quantify the height of the peaks $\nu_{\rm c}$ fixing the ratio of PBHs in dark matter $f_{\rm PBH}^{\rm tot}$. We consider the case of a monochromatic power spectrum $\mathcal{P}_{\zeta}=\mathcal{A}_{\zeta}\delta(\ln(k/k_{p}))$ as in section 2. To statistically compute the abundance of peaks leading to black hole formation, we follow the approach of [91, 92] based on the Gaussian statistics of $\zeta$ (see also [65] where the approach was also used in the context of the PTA analysis with arbitrary $w$).
@@
where $f(\nu)$ is the function introduced in Eq.(2.12). For the monochromatic spectrum, the number density of PBHs $n_{\rm PBH}(M)$ for a fixed mass $M$ is simply given by  [^6] Notice that for the general case, the relation between $n_{\rm PBH}(M)$ will involve integration over the variable $\xi$ introduced in section 2, but for the monochromatic power spectrum, this is simply given by $\xi=\nu$.
@@ Appendix C Suplemental figures of the gravitational collapse for $w=1/10$
In this appendix, we present additional figures illustrating the gravitational collapse for the case of $w=1/10$. Overall, the qualitative behaviour is similar to the cases discussed in section 4.1. We find that the most notable qualitative difference is that the collapse takes longer than in the $w=1/3$ case, as previously mentioned. Figs.16, 17 and 18 show a scenario where the fluctuation collapses to form a black hole. The initial shape of the fluctuation in the $z-x$ plane is nearly spherical, with eigenvalues $\lambda_{1}/(\nu\sigma_{2})=1.175$, $\lambda_{2}/(\nu\sigma_{2})=1.1$, and $\lambda_{3}/(\nu\sigma_{2})=0.825$, where $\lambda_{1}\approx\lambda_{2}$. In contrast, Figs.16, 17 and 18 depict a case with an initial prolate spheroidal geometry, where the eigenvalues are $\lambda_{1}/(\nu\sigma_{2})=0.825$, $\lambda_{2}/(\nu\sigma_{2})=1.35$, and $\lambda_{3}/(\nu\sigma_{2})=0.825$, with $\lambda_{1}=\lambda_{3}<\lambda_{2}$. Although the oscillating behaviour is similar to what was observed previously, it eventually transitions to an oblate geometry, approaching an almost spherical shape in the final stages.
Figure 16: Snapshots of the evolution of $\rho/\rho_{b}$ in the plane $y=0$ for $e=0.075$ and $p=-0.05$ with $w=1/10$.
Figure 17: Snapshots of the evolution of the Eulerian velocity $U^{x}$ in the plane $y=0$ for $e=0.075$ and $p=-0.05$ with $w=1/10$.
Figure 18: Snapshots of the evolution of the Eulerian velocity $U^{z}$ in the plane $y=0$ for $e=0.075$ and $p=-0.05$ with $w=1/10$.
Figure 19: Snapshots of the energy density $\rho/\rho_{b}$ (left-panel), lapse function $\alpha$ (middle-panel) and Eulerian velocity $U^{x}$ (right-panel) on the $x$ axis ($y=z=0$) for $e=0.08$ and $p=0$ with $w=1/10$.
Figure 20: Snapshots of the evolution of $\rho/\rho_{b}$ in the plane $y=0$ for $e=0$ and $p=-0.175$ with $w=1/10$.
Figure 21: Snapshots of the evolution of the Eulerian velocity $U^{x}$ in the plane $y=0$ for $e=0$ and $p=-0.175$ with $w=1/10$.
Figure 22: Snapshots of the evolution of the Eulerian velocity $U^{z}$ in the plane $y=0$ for $e=0,p=-0.175$ with $w=1/10$.
Figure 23: Snapshots of the energy density $\rho/\rho_{b}$ (left-panel), lapse function $\alpha$ (middle-panel) and Eulerian velocity $U^{x}$ (right-panel) on the $x$ axis ($y=z=0$) for $e=0$ and $p=0.175$ with $w=1/10$.

### v3

@@ 4.1 Dynamical evolution of the non-spherical gravitational collapse
We analyze two specific cases to examine the dynamics of gravitational collapse. In the first case, the fluctuation exceeds the critical threshold $\mu_{\rm t}>\mu_{\rm c}(e,p)]$, leading to black hole formation. In the second case, the fluctuation remains below the threshold $\mu_{\rm t}<\mu_{\rm c}(e,p)]$, resulting in the dispersion of the fluctuation. To study the dynamical evolution of the fluctuation, we plot the ratio of $\rho/\rho_{b}(t)$ and the 4-velocity of the fluid $U^{i}$. We focus on the case of radiation-dominated Universe $w=1/3$ and refer the reader to Appendix C for the $w=1/10$ case. We found the qualitative behaviour to be the same for both $w=1/3$ and $1/10$ except that the collapsing time is longer for $w=1/10$, which is consistent with the case of spherical simulations [29].
@@
In Figs. 4 and 5, we plot the velocity $U^{x}$ and $U^{z}$ on the $x$-$z$ plane, respectively, where we observe a highly non-spherical distribution. From the velocity plots of $U^{x}$ and $U^{z}$, we observe that the collapse initially progresses slightly faster along the $z$-axis compared to the $x$-axis, as indicated by the higher collapse velocity in $U^{z}$ (see the panels of $t/t_{H}\approx 0.04$). However, presumably because of the larger pressure gradient along the $z$-axis, the contraction along the $x$-axis overtakes the contraction along the $z$-axis (see the panels of $t/t_{H}\approx 40.04$ and compare the values of $U^{x}$ and $U^{z}$). Subsequently, the initial shape transitions from horizontally long to vertically long while preserving the ellipsoidal shape and reducing the size of the overdensity region (see the panels of $t/t_{H}=90.04$ and those after that in Fig. 3). At very late times, the shape becomes nearly spherical.
@@
Figure 6 shows the the energy density, lapse function, and fluid velocity in the $x$ components on the $x$-axis. Similar behaviors are found for $y$ and $z$ axes, and we do not display them. We observe a continuous decrease in the lapse function at the centre, and we infer the formation of an apparent horizon.
@@
Finally, Fig. 11 shows the time evolution of the lapse function at the centre $\alpha(\vec{r}=0)$ for different configurations. The figure highlights the detailed bouncing behaviour of the lapse function, which indicates apparent horizon formation when $\alpha(\vec{r}=0)$ continuously decreases and approaches to zero. We can observe that the time scale of the gravitational collapse is longer for a softer equation of state (smaller $w$). This can be understood as follows. In a simple model [85], the threshold of PBH formation is estimated by considering the critical situation in which the sound wave propagation time scale and the free-fall time scale are equal to each other. Since the sound wave speed is given by $\sqrt{w}$, the sound wave crossing time scale can be roughly estimated by $1/(\sqrt{w}H_{\rm ent})=3(1+w)/(2\sqrt{w})t_{\rm H}$. In the critical situation, this time scale coincides with the collapsing time scale. Therefore the gravitational collapse takes a much longer time with $w=1/10$ near the threshold  [^5] It should be noted that the values of the coordinate time shown in the figures are much larger than the estimated value because, with the gauge adopted in the numerical simulation, the time lapse is delayed in the central region.. Another effect to consider is pressure as a gravitational source: while they initially oppose the collapse, once the collapse is triggered, they primarily favor it by reducing the collapse time of the fluctuation, as they represent a form of gravitational energy. See, for instance, Fig. 1 of [29] for the case of spherical symmetry.
The top panels display cases with the fixed amplitude $\mu_{\rm t}$. Generally, we observe that non-spherical effects ($e,p\neq 0$) tend to slow down the collapse and can significantly increase the collapse time. The bottom panels show different configurations with varying amplitudes $\mu$ while fixing $p=0$. Here, we find that as $\mu$ increases beyond the critical value for spherical collapse, the deviation from sphericity required to prevent collapse also becomes larger, showing that non-spherical effects tend to prevent black hole formation. By running multiple simulations and analyzing the bouncing behaviour of the lapse function, we identified the threshold values, which will be discussed in detail in the next section.
@@ 4.3 Critical ellipticity as a function of the initial amplitude $\mu$
Figure 13: Critical value $e_{\rm c}$ with $p=0$ is depicted as a function of $(\mu-\mu_{\rm c,sp})/\mu_{\rm c,sp}$ together with the error bars given by the intervals between the parameter values of $e$ for the corresponding two cases of the black hole formation and dissipation. The dashed magenta lines correspond to the fitting to the equation $\tilde{e}_{c}(\mu)=\mathcal{K}_{e}[(\mu-\mu_{\rm c,sp})/\mu_{\rm c,sp}]^{\gamma_{e}}$, whose values $\mathcal{K}_{e}$ and $\gamma_{e}$ are shown in the plot for the case $w=1/10$ (blue) and $w=1/3$ (red).
In Fig. 13, we present the results of our simulations. Our numerical results closely follow a power-law behavior, described by,
@@
Our numerical results in Fig. 13 demonstrate that non-spherical effects make the gravitational collapse of fluctuations harder compared to the spherical case. Moreover, due to the behaviour of Eq.(4.2), even a slight deviation from sphericity can prevent a fluctuation from collapsing into a black hole if those fluctuations have an amplitude near the threshold in the critical regime. Let us evaluate, due to the non-spherical effects, how large fraction is excluded from the total number of black holes which is supposed to be formed without taking into account the non-sphericity. The remaining fraction can be estimated by integrating the probability distribution of $e$ and $p$ in the region enclosed by $p=\pm e$ and $p=\tilde{p}_{\rm c}(e)$ with a fixed value of $\mu$. For simplicity, we approximate the line $p=\tilde{p}_{\rm c}(e)$ by $e=\tilde{e}_{\rm c}(\mu)$ of a vertical line in the $e$-$p$ plane and integrate the probability distribution $\mathcal{P}_{e,p}$ over the region $\mathcal{R}$ defined by
@@
The integration of the probability distribution with the fitting function (4.2) can be found in Fig. 14 in percentatge $\%$”. For a typical value of $\mu$, which gives $(\mu-\mu_{\rm c,sp})/\mu_{\rm c,sp}\sim 1\%$, about $90\%$ of the configurations will collapse depending on $w$ and $\nu$ chosen, overcoming the non-spherical effects. This supports the conclusion reached in [18] for $w=1/3$ with an exponential Gaussian profile considering $p=\pm e$, which found that non-spherical effects do not significantly impact the threshold by more than $\sim 1\%$ [^6] Notice that, in [18], differently from this paper, another probability distribution, which is defined without considering the real space number density, is used for simplicity because this simplification does not cause any qualitative difference.. Our findings suggest that the total number of black holes can be accurately estimated based on the threshold given by spherically symmetric simulations for typical settings.
On the other hand, if we focus on the critical scaling regime, in which the amplitude is very close to $\mu_{\rm c,sp}$, the number of configurations collapsing into black holes can be significantly reduced. For instance, for $(\mu-\mu_{\rm c,sp})/\mu_{\rm c,sp}=10^{-3}$, most of the configurations are prevented from collapsing into black holes although it slightly depends on $w$ and $\nu$. Therefore the non-spherical effects can be highly significant in the critical regime, preventing a large fraction of configurations from collapsing into black holes. This behavior differs from the one when considering the Press-Schechter method [88, 89] (see, for instance, [90]), where all non-spherical configurations are considered equally probable, and therefore non-spherical effects are overestimated. Nevertheless, we report in the accompanying letter [91] that, in terms of the PBH mass function, the effects of the ellipticity may not have a significant impact even on the power-law small mass tail originates from the critical behavior.
Figure 14: Integration of the probability distribution of Eq. (2.11) integrated over the domain $\mathcal{R}$ defined in Eq.(4.3) in terms of $(\mu-\mu_{\rm c,sp})/\mu_{\rm c,sp}$ in percentatge value $\%$. Red and blue lines correspond to the cases with radiation and soft equations of state, respectively. The solid and dashed lines correspond to the cases for $\nu=9$ and $\nu=7$, respectively, which corresponds to putting the scale $k_{p}$ at different mass-scales; see Table1 and Appendix B for details. The two green vertical lines indicate $1\%$ and $2\%$, respectively.
@@
The fact that we do not observe a significant impact of non-spherical effects on gravitational collapse for $w=1/10$ compared to $w=1/3$ (as shown in Fig. 13) could be attributed to the following reason. First, we should note that the collapsing dynamics of the ellipsoidal system is different from the dust case and the well-known instability shown by Lin-Mestel-Shu [92] does not apply. Here let us try to understand our results following the Jeans criterion, which states that the system is unstable against gravitational collapse if the free-fall time scale is shorter than the sound wave crossing time scale. In our setting, at least for a relatively small ellipticity, the free-fall time scale $\sim 1/\sqrt{\rho}$ would be still relevant, and it does not change much due to the small ellipticity $e$ because the product $\lambda_{1}\lambda_{2}\lambda_{3}$ is conserved at the linear order of $e$ with $p=0$. On the other hand, the sound wave crossing time scale is expected to be shorter because the sound wave along the short axis may propagate through the system in a shorter time. The earlier bounce due to the shorter time scale may cause a larger circumference in the last stage of the collapse and prevent the black hole formation in terms of the hoop conjecture [93]. Therefore the gravitational collapse is expected to be harder. Although, with a given value of the initial amplitude $\mu$, this effect is expected to be more pronounced for larger $w$, the relative effect may be comparable since the threshold amplitude is smaller for a smaller value of $w$.
Once non-spherical effects become dominant impeding factors against the gravitational collapse as $w\rightarrow 0$, the Jeans criterion does not apply, and we need to consider different criteria. Non-sphericities are expected to grow during collapse [94, 92]. This could lead to complex dynamics associated with the deformation [95] and rotation of over-densities [32], and velocity dispersion potentially plays a significant role [96]. Further investigation is needed to clarify this aspect.
@@ 5 Conclusions
We have also examined how the threshold for black hole formation depends on the parameters $e$ and $p$. When we fixed the amplitude as a typical value $\mu_{\rm t}$, we found that the curve $p=p_{\rm c}(e)$, which describes the boundary of the region of black hole formation on the $e$-$p$ plane fits well a superellipse curve (described by Eq.(4.1)) characterized by an exponent $n^{\pm}$, where the sign $\pm$ denotes the two branches for the region $p\gtrless 0$. It would be interesting to investigate that functional form in terms of $\mu$, and its profile dependence for developing an analytical framework similar to spherically symmetric cases [28, 29] or including non-Gaussianities [67].
@@
Acknowledgments
A.E. acknowledges support from the JSPS Postdoctoral Fellowships for Research in Japan (Graduate School of Sciences, Nagoya University). C.Y. is supported in part by JSPS KAKENHI Grant Nos. 20H05850 and 20H05853.
@@ Appendix A Analytical formulas for the mean values of $e$ and $p$
In the large $\xi$ limit, the probability distribution Eq.(2.11) can be indeed approximated by a Gaussian and takes the simpler form,
@@ Appendix B High peaks for PBH formation assuming spherical symmetry
In this Appendix, under the assumption of spherical symmetry, we quantify the height of the peaks $\nu_{\rm c}$ fixing the ratio of PBHs in dark matter $f_{\rm PBH}^{\rm tot}$. We consider the case of a monochromatic power spectrum $\mathcal{P}_{\zeta}=\mathcal{A}_{\zeta}\delta(\ln(k/k_{p}))$ as in Sec. 2. To statistically compute the abundance of peaks leading to black hole formation, we follow the approach of [98, 99] based on the Gaussian statistics of $\zeta$ (see also [65] where the approach was also used in the context of the PTA analysis with arbitrary $w$).
@@
where $f(\nu)$ is the function introduced in Eq.(2.12). For the monochromatic spectrum, the number density of PBHs $n_{\rm PBH}(M)$ for a fixed mass $M$ is simply given by  [^7] Notice that for the general case, the relation between $n_{\rm PBH}(M)$ will involve integration over the variable $\xi$ introduced in Sec. 2, but for the monochromatic power spectrum, this is simply given by $\xi=\nu$.
@@ Appendix C Suplemental figures of the gravitational collapse for $w=1/10$
In this Appendix, we present additional figures illustrating the gravitational collapse for the case of $w=1/10$. Overall, the qualitative behaviour is similar to the cases discussed in Sec. 4.1. We find that the most notable qualitative difference is that the collapse takes longer than in the $w=1/3$ case, as previously mentioned. Figs. 16-19 show a scenario where the fluctuation collapses to form a black hole. The initial shape of the fluctuation in the zxz-x plane is nearly spherical, with eigenvalues $\lambda_{1}/(\nu\sigma_{2})=1.175$, $\lambda_{2}/(\nu\sigma_{2})=1.1$, and $\lambda_{3}/(\nu\sigma_{2})=0.825$, where $\lambda_{1}\approx\lambda_{2}$. In contrast, Figs. 20-23 depict a case with an initial prolate spheroidal geometry, where the eigenvalues are $\lambda_{1}/(\nu\sigma_{2})=0.825$, $\lambda_{2}/(\nu\sigma_{2})=1.35$, and $\lambda_{3}/(\nu\sigma_{2})=0.825$, with $\lambda_{1}=\lambda_{3}<\lambda_{2}$. Although the oscillating behaviour is similar to what was observed previously, it eventually transitions to an oblate geometry, approaching an almost spherical shape in the final stages.
Figure 16: Snapshots of the evolution of $\rho/\rho_{b}$ in the plane $y=0$ for $e=0.075$ and p=0.05p=-0.05 with $w=1/10$.
Figure 17: Snapshots of the evolution of the Eulerian velocity $U^{x}$ in the plane $y=0$ for $e=0.075$ and p=0.05p=-0.05 with $w=1/10$.
Figure 18: Snapshots of the evolution of the Eulerian velocity $U^{z}$ in the plane $y=0$ for $e=0.075$ and p=0.05p=-0.05 with $w=1/10$.
Figure 19: Snapshots of the energy density $\rho/\rho_{b}$ (left-panel), lapse function $\alpha$ (middle-panel) and Eulerian velocity $U^{x}$ (right-panel) on the $x$ axis ($y=z=0$) for $e=0.075$ and p=0.05p=-0.05 with $w=1/10$.
Figure 20: Snapshots of the evolution of $\rho/\rho_{b}$ in the plane $y=0$ for $e=0$ and p=0.175p=-0.175 with $w=1/10$.
Figure 21: Snapshots of the evolution of the Eulerian velocity $U^{x}$ in the plane $y=0$ for $e=0$ and p=0.175p=-0.175 with $w=1/10$.
Figure 22: Snapshots of the evolution of the Eulerian velocity $U^{z}$ in the plane $y=0$ for e=0,p=0.175e=0,p=-0.175 with $w=1/10$.
Figure 23: Snapshots of the energy density $\rho/\rho_{b}$ (left-panel), lapse function $\alpha$ (middle-panel) and Eulerian velocity $U^{x}$ (right-panel) on the $x$ axis ($y=z=0$) for $e=0$ and p=0.175p=-0.175 with $w=1/10$.